OpenStudy (anonymous):

solve the given equation in term of 2kpi tan^2(x)-2sec(x)=2

myininaya (myininaya):

We need to put this in terms of one trig function Recall: \[1+\tan^2(x)=\sec^2(x)\] => \[\tan^2(x)=\sec^2(x)-1\] So we have the equation \[\tan^2(x)-2\sec(x)=2\] To put in terms of sec(x) we will use the identity above \[(\sec^2(x)-1)-2\sec(x)=2\] So we have \[\sec^2(x)-1-2\sec(x)=2\] Now subtracting 2 on both sides and reordering the terms we have \[\sec^2(x)-2\sec(x)-1-2=0\] Combining like terms gives \[\sec^2(x)-2\sec(x)-3=0\] Now do you know how to solve: \[u^2-2u-3=0\] If you know how to solve that equation then you will be able to solve the first since these equations have the relationship \[u=\sec(x)\] So tell we what you get when you solve \[u^2-2u-3=0\] And then we will go from there.

OpenStudy (anonymous):

are you using calculus to solve this?

myininaya (myininaya):

No

myininaya (myininaya):

I used a trig identity

OpenStudy (anonymous):

so....(secx-2)=0 (secx+1)=0???

myininaya (myininaya):

No

myininaya (myininaya):

\[u^2-2u-3=0\] Try factoring this first You will get (sec(x)-3)(sec(x)+1)=0

myininaya (myininaya):

since -3(1)=-3 and -3+1=-2

myininaya (myininaya):

Now set both factors =0

OpenStudy (anonymous):

secx=3 secx=-1

myininaya (myininaya):

Right

OpenStudy (anonymous):

the answer in the book doesnt make sense

myininaya (myininaya):

Why do you say that?

OpenStudy (anonymous):

because its in decimal form, I'm confused on when to use exact form and decimal. or does that not even matter?

myininaya (myininaya):

It does matter What does the question say? Does it say you can round?

myininaya (myininaya):

Usually it says what to do

myininaya (myininaya):

It shouldn't be a guessing name If it doesn't say anything about rounding then you provide the exact answer

OpenStudy (anonymous):

solve the given equation

myininaya (myininaya):

gosh game*

myininaya (myininaya):

sorry

OpenStudy (anonymous):

in the book the answers for that area jump from decimal to exact

myininaya (myininaya):

ok if it doesn't say anything about rounding i would put the exact answer regardless of what the back of the book says

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answers are: (2k+1)pi, 1.23+2kpi, 5.05+2kpi

myininaya (myininaya):

Those are approximations though

myininaya (myininaya):

well the last two

OpenStudy (anonymous):

Also confused on how they got to that point

myininaya (myininaya):

ok so we have sec(x)=3 or sec(x)=-1

OpenStudy (anonymous):

ok

myininaya (myininaya):

Or we could say since \[\sec(x+2 k \pi) = \sec(x)\] then \[\sec(x+ 2 k \pi)=3 \text{ or } \sec(x+2 k \pi ) =-1\]

myininaya (myininaya):

Now take arcsec( ) of both sides

myininaya (myininaya):

\[=> x+2 k \pi = \sec^{-1}(3) \text{ or } x+2 k \pi =\sec^{-1}(-1)\]

OpenStudy (anonymous):

oh is that where the approximations come from?

myininaya (myininaya):

yep i assume they approximate arcsec(3) but arcsec(-1) you can evaluate that exactly

myininaya (myininaya):

But....

OpenStudy (anonymous):

ok now I think I understand.

myininaya (myininaya):

Since cos(x) is an even function then sec(x) is an even function We do have more solutions

myininaya (myininaya):

\[\sec(x)=\sec(-x)\]

myininaya (myininaya):

So we have the above+ \[x+2k \pi=-\sec^{-1}(3) \text{ or } x+2 k \pi=-\sec^{-1}(-1)\]

OpenStudy (anonymous):

so problems like this i would break down the first part of the equation to match the second part and take the arc of the identity?

myininaya (myininaya):

Sometimes and also you might need to use the fact that a trig function is odd or even like i did

OpenStudy (anonymous):

Ok. I'm going to try a few problems thanks

myininaya (myininaya):

Np So I gave you 4 equations that will give you your solution just so you know :)